# Kähler compactification of $\mathbb{C}^{n}$ and Reeb dynamics

### v1

@@ Abstract.
Let $X$ be a smooth complex manifold. Assume that $Y\subset X$ is a Kähler submanifold such that $X\setminus Y$ is biholomorphic to $\mathbb{C}^{n}$. We prove that $(X,Y)$ is biholomorphic to $(\mathbb{P}^{n},\mathbb{P}^{n-1})$. We then study certain Kähler orbifold compactifications of $\mathbb{C}^{n}$ and, as an application, prove that on $\mathbb{C}^{3}$ the flat metric is the only asymptotically conical Ricci-flat Kähler metric whose metric cone at infinity has a smooth link. As a key technical ingredient, a new formula for minimal discrepancy of isolated Fano cone singularities in terms of Conley-Zehnder indices in symplectic geometry is derived.
@@ 1. Introduction
Our work is motivated by two problems in complex geometry. The first is Hirzebruch’s classical problem on analytic compactification of $\mathbb{C}^{n}$. Let $X$ be a compact complex manifold of complex dimension $n$. Assume that $Y$ is a subvariety of $X$ such that $X\setminus Y$ is biholomorphic to $\mathbb{C}^{n}$. Of course, the simplest example is $(X,Y)=(\mathbb{P}^{n},\mathbb{P}^{n-1})$ where $\mathbb{P}^{n-1}\subset\mathbb{P}^{n}$ is a linear subspace. In this case, we will just say that $(X,Y)$ is standard. In [Hir54], Hirzebruch asked the problem of classifying all such pair $(X,Y)$ with the second Betti number $1$. When $\dim X=1$ or $2$, it is known that there is only the standard example. For $n=3$, under the assumption that $X$ is projective, all such analytic compactifications of $\mathbb{C}^{3}$ are now classified. The subvariety $Y$ in any known non-standard example is always singular. For example, one can compactify $\mathbb{C}^{n}$ to a smooth quadric hypersurface in $\mathbb{P}^{n+1}$ by adding an irreducible singular divisor (a projective cone over a lower dimensional quadric) at infinity. It is a well-known conjecture that if $Y$ is assumed to be smooth then $(X,Y)$ should be standard (see [BM78], [PS91]). This was proved in [BM78] when $\dim X\leq 3$. Here we prove this conjecture under the Kähler assumption for all dimensions.
@@
With the above notation, if $Y\subset X$ is a Kähler submanifold, $(X,Y)$ is standard.
Early results of Brenton-Morrow from [BM78] showed already that, with the same Kähler assumption, $Y$ must be a Fano manifold and they reduce the problem to proving that $Y$ is biholomorphic to $\mathbb{P}^{n-1}$. We will prove that the Fano index of $Y$ must be $n$ which then implies that $Y$ is $\mathbb{P}^{n-1}$ by a well-known criterion of Kobayashi-Ochiai ([KO73]). The main tools we use are ($S^{1}$-equivariant, positive) symplectic homology and two ways to compute it, one from the functorial properties and the special symplectic topology of $\mathbb{C}^{n}$, and the other from the Morse-Bott spectral sequences applied to the Boothy-Wang contact structure on the unit circle bundle of the normal bundle of $Y$.
Another problem that motivates our work is the classification of asymptotically conical (AC) Ricci-flat metrics, which has been recently studied extensively. Let $(W,g)$ be a complete Kähler manifold. It is called asymptotically conical (AC) if it is asymptotical to a Kähler cone $(\mathcal{C},g_{0})$ (see Definition 2.3). By the compactification results from [CH24, Li20] (see Theorem 1.3), the classification of such Kähler metrics are closely related to the orbifold compactifications of $W$. We have the following folklore conjecture  [^1] A version of such kind of conjecture was proposed by G. Tian. :
@@
Note that by the recent examples constructed in [Li19, CR21, Sze19], the condition of having a smooth link is needed. We will use the symplectic techniques in an orbifold setting and reduce the conjecture 1.2 to Shokurov’s conjecture (see Conjecture 2.1) for isolated cone singularities.
@@
    Let $\mathcal{Y}\subset\mathcal{X}$ be an orbifold Kähler compactification of $\mathbb{C}^{n}$ that is associated to a complete AC Kähler metric. Then the orbifold cone $\mathcal{C}(\mathcal{Y},\mathcal{L})$ over $\mathcal{Y}$ with $\mathcal{L}$ being the orbifold normal bundle is a Gorenstein klt singularity which has the minimal discrepancy equal to $n-1$.
@@
    Conjecture 1.2 is true assuming Shokurov’s conjecture. In particular, Conjecture 1.2 is true when $n=3$.
As a key ingredient for proving the above results, we derive the following new formula for the minimal discrepancy of isolated Fano cone singularities (see 2.1 for the explanation of these terminologies) in terms of the $\mathrm{lSFT}$-invariant (see Definition 2.12), which was introduced by McLean ([McL16]) as a generalized Conley-Zehnder type index for closed Reeb orbits.
@@
(1) $22\mathrm{md}(o,\mathcal{C})=\inf_{\gamma}\mathrm{lSFT}_{\eta}(\gamma)>-$ 2 .
@@
(2) $32\mathrm{md}(o,\mathcal{C})=\inf\{d\;|\;SH^{+,S^{1}}_{d}(W;\mathbb{Q})\neq 0\}+n-$ 3
@@
In the more special case of isolated quotient singularities, this formula recovers the result in [Zho24, Proposition 2.12]. We remark that our proof of Theorem 1.4 does not use the main formula from [McL16, Theorem 1.1] which states that $2\mathrm{md}$ is equal to the highest minimal index denoted by $\mathrm{hmi}$ for canonical singularities. In fact, the definition of the contact invariant $\mathrm{hmi}$ in [McL16] requires enumeration through all contact forms, while 1.4 asserts that the minimal discrepancy can be computed using the Reeb dynamics from a single conic contact form and is witnessed by a non-trivial class in Floer homology. The latter can be approached using properties of symplectic homology, which allow us to show the equality $\mathrm{md}(o,\mathcal{C})=n-1$ for the Fano cone arising in our problems. The existence of a Liouville filling $W$ is just for the existence of $\mathbb{Q}$-graded $S^{1}$-equivariant positive symplectic cohomology. Such a Liouville filling always exists if $o$ is smoothable and $\dim_{\mathbb{C}}\mathcal{C}\geq 3$. The specific choice of $W$ does not matter. In a more general setting, one can replace $SH^{+,S^{1}}_{*}(W;\mathbb{Q})$ by the linearized contact homology of $M$ w.r.t. a $\mathbb{Q}$-graded augmentation over $\mathbb{C}$, which always exists in this case, see 3.8 for more discussions.
The application of tools from symplectic geometry to compactification problem seems new to us. From a broader perspective, we use techniques from symplectic geometry to rule out certain singularities. The application of (very different) symplectic techniques to complex geometry to rule out singularities has appeared in the recent solution to the generalized Bishop problem [FHYZ24].
@@ Acknowledgments
C. Li is partially supported by NSF (Grant No. DMS-2305296). Z. Zhou is supported by National Key R&D Program of China under Grant No.2023YFA1010500, NSFC-12288201 and NSFC-12231010. C. Li would like to thank Prof. Laszlo Lempert for helpful discussions, and Prof. Mircea Mustaţă for comments on Shokurov’s conjecture. Both authors would like to thank Prof. Xiaojun Huang for putting us into connection.
@@ 2.1. Fano cone singularities and AC Kähler metrics
Let $\mathcal{C}$ be a normal affine variety with log terminal $\mathbb{Q}$-Gorenstein singularities. Here $\mathbb{Q}$-Gorenstein means that the canonical divisor $K_{\mathcal{C}}$ is $\mathbb{Q}$-Cartier. To define the log terminal condition, we choose a log resolution of singularities $\mu:\tilde{\mathcal{C}}\rightarrow\mathcal{C}$ such that the inverse image of the singular locus of $\mathcal{C}$ is given by a simple normal crossing divisor $\bigcup_{i}E_{i}$. We can then write an identity:
$K_{\tilde{\mathcal{C}}}=\mu^{*}K_{\mathcal{C}}+\sum_{i}a(E_{i},\mathcal{C})E_{i}.$
The affine variety $\mathcal{C}$ has log terminal singularities if $a(E_{i},\mathcal{C})>-1$ for any $i$. Log terminal singularities are local analogues of Fano varieties and they play an important role in birational algebraic geometry. From now on, we always assume that $o\in\mathcal{C}$ is an isolated singularity. So we can choose a resolution of singularity $\mu:\tilde{\mathcal{C}}\rightarrow\mathcal{C}$ that is an isomorphism in a punctured neighborhood of $o\in\mathcal{C}$ such that $\mu^{-1}(o)$ is a simple normal crossing divisor. The minimal discrepancy of the klt singularity $o\in\mathcal{C}$ is then defined as $\mathrm{md}(o,\mathcal{C})=\min_{i}\{a(E_{i},\mathcal{C});\mu(E_{i})=o\}$. One can show that this quantity does not depend on the choice of log resolutions. We have the following important conjecture:
@@
Shokurov’s conjecture has been proved in dimension $\leq 3$ ([Mar96, Kaw93]).
We now explain the concept of a Kähler cone metric on $\mathcal{C}$. This means that there is smooth function $\mathfrak{r}:\mathcal{C}^{\circ}:=\mathcal{C}\setminus o\rightarrow\mathbb{R}_{>0}$, called the radius function, which is a surjective submersion such that if we set $\hat{\omega}:=\sqrt{-1}\partial\bar{\partial}{\mathfrak{r}}^{2}$ and $g_{0}=\hat{\omega}(\cdot,J_{0}\cdot)$ where $J_{0}$ is the integrable complex structure on $\mathcal{C}^{\circ}$, then $g_{0}$ is a Riemannian metric on $\mathcal{C}^{\circ}$ that is isometric to a Riemannian cone metric $d{\mathfrak{r}}^{2}+{\mathfrak{r}}^{2}g^{M}$ where $M=\{{\mathfrak{r}}=1\}$ is the link and $g^{M}=g_{0}|_{M}$. Define the conic contact form associated to the radius function ${\mathfrak{r}}$ as $\eta=-{\mathfrak{r}}^{-1}J_{0}d{\mathfrak{r}}$. Then we have $\hat{\omega}=2{\mathfrak{r}}d{\mathfrak{r}}\wedge\eta+{\mathfrak{r}}^{2}d\eta$. The Reeb vector field, denoted by ${\mathfrak{v}}=J_{0}(\mathfrak{r}\partial_{\mathfrak{r}})$, is a holomorphic Killing vector field on $\mathcal{C}$ and satisfies $\eta({\mathfrak{v}})=1$. The associated holomorphic vector field ${\mathfrak{v}}_{\mathbb{C}}={\mathfrak{r}}\partial_{\mathfrak{r}}-\sqrt{-1}J_{0}({\mathfrak{r}}\partial_{\mathfrak{r}})$ generates an effective holomorphic action by a torus $(\mathbb{C}^{*})^{m}$ which we will also denote by $\langle{\mathfrak{v}}\rangle$. If $m=1$, ${\mathfrak{v}}$ and $\eta$ are called quasi-regular and otherwise (i.e. when $m\geq 2$) ${\mathfrak{v}}$ and $\eta$ are called irregular.
@@
In the quasi-regular case, we can take the quotient of $\mathcal{C}^{\circ}$ by the $\mathbb{C}^{*}$-action or equivalently the quotient of $M$ by the $S^{1}$-action to get an orbifold $\mathcal{Y}:=(Y,\Delta)=M/S^{1}=\mathcal{C}^{\circ}/\mathbb{C}^{*}$ equipped with an orbifold line bundle $\mathcal{L}$. There is a natural projection $\pi:\mathcal{C}^{\circ}\rightarrow Y$ which makes $\mathcal{C}^{\circ}$ a $\mathbb{C}^{*}$-Seifert bundle over the cyclic orbifold $\mathcal{Y}$ which was studied for example by Kollár in [Kol04]. The affine variety $\mathcal{C}$ is algebraically described as:
@@
Geometrically it is obtained from the total space of the orbifold line bundle $\mathcal{L}^{-1}$ by contracting the zero section. For simplicity, We will call $o\in\mathcal{C}$ an orbifold cone singularity. $\mathcal{C}$ can be compactified to a projective variety $\overline{\mathcal{C}}$ by adding an divisor $\mathcal{Y}_{\infty}\cong\mathcal{Y}$ at infinity (see [Kol04, 14]). According to [Kol04, 42], the singularity $o\in\mathcal{C}$ is log terminal if and only $(Y,\Delta)$ is a Fano orbifold (i.e. the orbifold anti-canonical line bundle $-K^{\mathrm{orb}}_{(Y,\Delta)}$ is ample) and there exists $\ell\in\mathbb{Q}_{>0}$ such that $-K^{\mathrm{orb}}_{(Y,\Delta)}\sim_{\mathbb{Q}}\ell\mathcal{L}$ (see Proposition 2.10 for a proof in our setting). In this case, we will simply call the orbifold cone $o\in\mathcal{C}$ a Fano cone (following the terminology in [LLX20, Definition 3.7]). Moreover, in the quasi-regular case, the metrics $g_{0}$ and $g^{M}$ discussed in the previous paragraph are induced by a Hermitian metric $h$ on $\mathcal{L}$ over $(Y,\Delta)$ whose Chern curvature will be denoted by $\omega^{Y}$.
@@
Let $(W,g)$ be a complete Kähler manifold whose integrable complex structure is denoted by $J$. It is called asymptotically conical (AC) with the asymptotical cone $(\mathcal{C},g_{0})$ if there exists a compact subset $K\subset W$ and a diffeomorphism $\Phi:\{{\mathfrak{r}}>1\}\rightarrow M\setminus K$ such that for some $\lambda_{1}>0,\lambda_{2}>0$ and all $j\in\mathbb{Z}_{\geq 0}$.
@@
We will use the following compactification result which only assumes that the complex structure is asymptotical to the cone.
@@
Let $(W,g,J)$ be a complete Kähler manifolds and $(\mathcal{C},g_{0},J_{0})$ be an orbifold cone as before. Assume that there exist a compact subset $K\subset W$ and a diffeomorphism $\Phi:\{\mathfrak{r}>1\}\rightarrow M\setminus K$ such that for some $\lambda_{2}>0$ and all $j\in\mathbb{Z}_{\geq 0}$,
@@
    ([Li20, CH24]) Assume that the Reeb vector field ${\mathfrak{v}}$ is quasi-regular on the cone $\mathcal{C}$. There exists a compact Kähler orbifold $\mathcal{X}$ and a sub-orbifold divisor $\mathcal{Y}$ such that $\mathcal{Y}\cong\mathcal{C}^{\circ}/\langle\mathfrak{v}\rangle$ and $\mathcal{X}\setminus\mathcal{Y}$ is biholomorphic to $W$. Moreover there exists an orbifold diffeomorphism $\Phi:(\mathcal{U},\mathcal{Y})\rightarrow(\mathcal{V},\mathcal{Y}_{\infty})$ where $\mathcal{U}$ is a neighborhood of $\mathcal{Y}\subset\mathcal{X}$ and $\mathcal{V}$ is a neighborhood of $\mathcal{Y}_{\infty}\subset\overline{\mathcal{C}}$.

### v3

@@ Abstract.
Let $X$ be a smooth complex manifold. Assume that $Y\subset X$ is a Kähler submanifold such that $X\setminus Y$ is biholomorphic to $\mathbb{C}^{n}$. We prove that $(X,Y)$ is biholomorphic to $(\mathbb{P}^{n},\mathbb{P}^{n-1})$. We then study certain Kähler orbifold compactifications of $\mathbb{C}^{n}$ and, as an application, prove that on $\mathbb{C}^{3}$ the flat metric is the only asymptotically conical Ricci-flat Kähler metric whose metric cone at infinity has a smooth link. As a key technical ingredient, we derive a new characterization of minimal discrepancy of isolated Fano cone singularities by using $S^{1}$-equivariant positive symplectic homology.
@@ 1. Introduction
Our work is motivated by two problems in complex geometry. The first is Hirzebruch’s classical problem on analytic compactification of $\mathbb{C}^{n}$. Let $X$ be a compact complex manifold of complex dimension $n$. Assume that $Y$ is a subvariety of $X$ such that $X\setminus Y$ is biholomorphic to $\mathbb{C}^{n}$. Of course, the simplest example is $(X,Y)=(\mathbb{P}^{n},\mathbb{P}^{n-1})$ where $\mathbb{P}^{n-1}\subset\mathbb{P}^{n}$ is a linear subspace. In this case, we will just say that $(X,Y)$ is standard. In [Hir54], Hirzebruch asked the problem of classifying all such pairs $(X,Y)$ with the second Betti number $b_{2}(X)=1$. The Betti number condition is known to imply that $Y$ is irreducible. When $\dim X=1$ or $2$, it is known that there is only the standard example. For $n=3$, under the assumption that $X$ is projective, all such analytic compactifications of $\mathbb{C}^{3}$ are now classified. The subvariety $Y$ in any known non-standard example is always singular. For example, one can compactify $\mathbb{C}^{n}$ to a smooth quadric hypersurface in $\mathbb{P}^{n+1}$ by adding an irreducible singular divisor (which is a projective cone over a lower-dimensional quadric) at infinity. It is a long-standing conjecture that if $Y$ is assumed to be smooth then $(X,Y)$ should be standard (see [vdV62], [BM78], [PS91]). This was proved in [BM78] when $\dim X\leq 3$. Here we prove this conjecture under the Kähler assumption for all dimensions.
@@
With the above notation, if $Y$ is a Kähler submanifold, $(X,Y)$ is standard.
Early results of Brenton-Morrow from [BM78] showed already that, with the Kähler assumption, $Y$ must be a Fano manifold and they reduce the problem to proving that $Y$ is biholomorphic to $\mathbb{P}^{n-1}$ (see Theorem 2.8). We will prove that the Fano index of $Y$ must be $n$, which then implies that $Y$ is $\mathbb{P}^{n-1}$ by a well-known criterion of Kobayashi-Ochiai ([KO73]). The main tools we use are ($S^{1}$-equivariant, positive) symplectic homology and two ways to compute it, one from the functorial properties and the special symplectic topology of $\mathbb{C}^{n}$, and the other from the Morse-Bott spectral sequences applied to the Boothby-Wang contact structure on the unit circle bundle of the normal bundle of $Y$.
Another problem that motivates our work is the study of asymptotically conical (AC) Ricci-flat Kähler metrics, which goes back to the work of Tian-Yau [TY91]. Let $(W,g)$ be a complete Kähler manifold. It is called asymptotically conical (AC) if it is asymptotic to a Kähler cone $(\mathcal{C},g_{0})$ (see Definition 2.5). By the compactification results from [CH24, Li20] (see Theorem 2.7), the classification of such Kähler metrics is closely related to the orbifold compactifications of $W$. We have the following folklore conjecture:
@@
In [Tia06, Theorem 5.2], Tian proved that this is true when $n=2$, answering a question of Calabi, and in this case, the condition of the asymptotical metric cone $\mathcal{C}$ having a smooth link automatically holds true thanks to a regularity result of Cheeger-Tian. He then asked a similar question in higher dimensions (see [Tia06, Remark 5.3]). By the recent examples constructed in [Li19, CR21, Sze19], when $n\geq 3$, the condition of having a smooth link is needed. We will use the symplectic techniques in an orbifold setting and reduce the Conjecture 1.2 to Shokurov’s conjecture (see Conjecture 2.3) for isolated cone singularities.
@@
    Let $\mathcal{Y}\subset\mathcal{X}$ be an orbifold Kähler compactification of $\mathbb{C}^{n}$ that is associated to a complete AC Kähler metric (obtained from Theorem 2.7). Then the orbifold cone $\mathcal{C}(\mathcal{Y},\mathcal{L})$ over $\mathcal{Y}$ with $\mathcal{L}$ being the orbifold normal bundle is a Gorenstein klt singularity which has the minimal discrepancy equal to n1n-1.
@@
    Conjecture 1.2 is true assuming Shokurov’s conjecture. In particular, Conjecture 1.2 is true when $n\leq 3$.
As a key ingredient for proving the above results, we derive the following new formula for the minimal discrepancy of isolated Fano cone singularities (see §2.1 for the explanation of these terminologies). Our formula is motivated by the work of McLean ([McL16]), who introduced the $\mathrm{lSFT}$-invariant (see Definition 2.18) as a generalized Conley-Zehnder type index for closed Reeb orbits.
@@
(1) $2\,\mathrm{md}(o,\mathcal{C})=\inf_{\gamma}\mathrm{lSFT}_{\eta}(\gamma)>-2.$
@@
(2) $2\,\mathrm{md}(o,\mathcal{C})=\inf\{d\;|\;SH^{+,S^{1}}_{d}(W;\mathbb{Q})\neq 0\}+n-3$
@@
In the more special case of isolated quotient singularities, this formula recovers the result in [Zho24, Proposition 2.12]. We remark that our proof of Theorem 1.4 does not use the main formula from [McL16, Theorem 1.1] (and in fact reproves it in Proposition 3.8 for our setting) which states that $2\cdot\mathrm{md}$ is equal to the highest minimal index denoted by $\mathrm{hmi}$ for canonical singularities. Indeed, the definition of the contact invariant $\mathrm{hmi}$ in [McL16] (see (18)) requires enumeration through all contact forms [^1] The minimal discrepancy is achieved by certain contact forms constructed from a simple normal crossing resolution in [McL16], while 1.4 asserts that the minimal discrepancy can be computed using the Reeb dynamics from a single conic contact form and is witnessed by a non-trivial class in Floer homology. The latter can be approached using properties of symplectic homology, which allows us to show the equality $\mathrm{md}(o,\mathcal{C})=n-1$ for the Fano cone arising in our problems. The existence of a Liouville filling $W$ is just for the existence of $\mathbb{Q}$-graded $S^{1}$-equivariant positive symplectic cohomology. The specific choice of $W$ does not matter. In a more general setting, one can replace $SH^{+,S^{1}}_{*}(W;\mathbb{Q})$ by the linearized contact homology of $M$ w.r.t. a $\mathbb{Q}$-graded augmentation over $\mathbb{C}$, which always exists in our case (see 3.9 for more discussions).
The application of tools from symplectic geometry to the compactification problem seems new to us. From a broader perspective, we use techniques from symplectic geometry to rule out certain singularities. The application of (very different) symplectic techniques to complex geometry to rule out singularities has appeared in the recent solution to the generalized Bishop problem [FHYZ24].
@@ Acknowledgments
C. Li is partially supported by NSF (Grant No. DMS-2305296). Z. Zhou is supported by National Key R&D Program of China under Grant No.2023YFA1010500, NSFC-12288201, and NSFC-12231010. C. Li would like to thank Prof. Laszlo Lempert for helpful discussions, Prof. Mircea Mustaţă for comments on Shokurov’s conjecture, and Prof. Gang Tian for his interest and useful comments. Both authors would like to thank Prof. Xiaojun Huang for putting them in contact. They also thank Prof. Thomas Peternell for communications regarding his preprint [Pet24].
@@ 2.1. Fano cone singularities and AC Kähler metrics
Let $\mathcal{C}$ be an $n$-dimensional normal affine variety with an isolated $\mathbb{Q}$-Gorenstein singularity $o\in\mathcal{C}$. Here $\mathbb{Q}$-Gorenstein means that the canonical divisor $K_{\mathcal{C}}$ is $\mathbb{Q}$-Cartier, i.e. there exists $k\in\mathbb{Z}_{>0}$ such that $kK_{\mathcal{C}}$ is a Cartier divisor. We choose a log resolution of singularity $\mu:\tilde{\mathcal{C}}\rightarrow\mathcal{C}$ such that $\mu$ is an isomorphism away from $o\in\mathcal{C}$ and $\mu^{-1}(o)$ is a simple normal crossing divisor. Then we have an identity
(3) $K_{\tilde{\mathcal{C}}}=\mu^{*}K_{\mathcal{C}}+\sum_{i}a(E_{i},\mathcal{C})E_{i}$
where $a(E_{i},\mathcal{C})\in\mathbb{Q}$ is called the discrepancy of the exceptional divisor $E_{i}$.
Remark 2.1.
Analytically, one can compute discrepancies in the following way. For simplicity, we assume $k=1$ (Gorenstein case) and refer to [Kol13, Definition 2.1] for the general situation. Pick a local generator $s$ of $\mathcal{O}(K_{\mathcal{C}})(U)$ where $U$ is a small neighborhood of $o\in\mathcal{C}$. Then $s$ is a non-vanishing holomorphic $n$-form on $U\setminus o$ and its pull-back $\mu^{*}s$ is in general a meromorphic $n$-form. Choose a general point on $E_{i}$ and nearby coordinates $\{y_{1},\dots,y_{n}\}$ such that $E_{i}=\{y_{1}=0\}$ and $\mu^{*}s=f(y)dy_{1}\wedge\cdots\wedge dy_{n}$. Then by definition $a(E_{i},\mathcal{C})=\mathrm{ord}_{\{y_{1}=0\}}f(y)$. For example, if $E$ is the exceptional divisor of the standard blowup of $\mathbb{C}^{n}$ at a point, then $a(E,\mathbb{C}^{n})=n-1$ .
Definition 2.2.
With the above notation, the singularity $(\mathcal{C},o)$ is Kawamata log terminal (klt) if $a(E_{i},\mathcal{C})>-1$ for any $i$. The minimal discrepancy of $o\in\mathcal{C}$ is defined as
$\mathrm{md}(o,\mathcal{C})=\min_{i}\left\{a(E_{i},\mathcal{C});\mu(E_{i})=\{o\}\right\}.$
In other words, the singularity $(\mathcal{C},o)$ is klt if and only if $\mathrm{md}(o,\mathcal{C})>-1$. It is known that this definition of klt singularity and the minimal discrepancy do not depend on the choice of log resolution of the singularity. Klt singularities are local analogues of Fano varieties, and they play an important role in birational algebraic geometry (see [Kol13]).
@@
Shokurov’s conjecture has important applications in birational algebraic geometry, and it has been proved in dimension $\leq 3$ ([Mar96, Kaw93]).
We now explain the concept of a Kähler cone metric on $\mathcal{C}$. This means that there is smooth function $\mathfrak{r}:\mathcal{C}^{\circ}:=\mathcal{C}\setminus o\rightarrow\mathbb{R}_{>0}$, called the radius function, which is a surjective submersion such that if we set $\hat{\omega}:=\sqrt{-1}\partial\bar{\partial}{\mathfrak{r}}^{2}$ and $g_{0}=\hat{\omega}(\cdot,J_{0}\cdot)$ where $J_{0}$ is the integrable complex structure on $\mathcal{C}^{\circ}$, then $g_{0}$ is a Riemannian metric on $\mathcal{C}^{\circ}$ that is isometric to a Riemannian cone metric $d{\mathfrak{r}}^{2}+{\mathfrak{r}}^{2}g^{M}$ where $M=\{{\mathfrak{r}}=1\}$ is the link and $g^{M}=g_{0}|_{M}$. Define the conic contact form associated to the radius function $\mathfrak{r}$} as $\eta=-{\mathfrak{r}}^{-1}J_{0}d{\mathfrak{r}}$. Then we have $\hat{\omega}=2{\mathfrak{r}}d{\mathfrak{r}}\wedge\eta+{\mathfrak{r}}^{2}d\eta$. The Reeb vector field, denoted by ${\mathfrak{v}}=J_{0}(\mathfrak{r}\partial_{\mathfrak{r}})$, is a holomorphic Killing vector field on $\mathcal{C}$ and satisfies $\eta({\mathfrak{v}})=1$. The associated holomorphic vector field ${\mathfrak{v}}_{\mathbb{C}}={\mathfrak{r}}\partial_{\mathfrak{r}}-\sqrt{-1}J_{0}({\mathfrak{r}}\partial_{\mathfrak{r}})$ generates an effective holomorphic action by a (complexified) torus $(\mathbb{C}^{*})^{m}$ which we will also denote by $\langle{\mathfrak{v}}\rangle$. If $m=1$, $\mathfrak{v}$} and $\eta$ are called quasi-regular and otherwise (i.e. when $m\geq 2$) $\mathfrak{v}$} and $\eta$ are called irregular.
[…]
@@
In the quasi-regular case, we can take the quotient of $\mathcal{C}^{\circ}$ by the $\mathbb{C}^{*}$-action or, equivalently, the quotient of $M$ by the $S^{1}$-action, to get an orbifold $\mathcal{Y}:=(Y,\Delta)=M/S^{1}=\mathcal{C}^{\circ}/\mathbb{C}^{*}$ equipped with an orbifold line bundle $\mathcal{L}$. Here $Y$ denotes the quotient space and $\Delta=\sum_{\alpha}(1-\frac{1}{m_{\alpha}})D_{\alpha}$ denotes the branch divisor where the sum is taken over all Weil divisors $D_{\alpha}$ that lie in the orbifold singular locus and $m_{\alpha}$ is the gcd of the orders of the local uniformizing groups taken over all points of $D_{\alpha}$ (see [BG08, Definition 4.4.8]). There is a natural projection $\pi:\mathcal{C}^{\circ}\rightarrow Y$ which makes $\mathcal{C}^{\circ}$ a $\mathbb{C}^{*}$-Seifert bundle over the cyclic orbifold $\mathcal{Y}$ which was studied by Kollár in [Kol04] [^2] In the literature of algebraic geometry, $\mathcal{C}$ is also called an affine variety with a good $\mathbb{C}^{*}$-action (See [Kol04] and its references). . The affine variety $\mathcal{C}$ is algebraically described as:
@@
Geometrically, it is obtained from the total space of the orbifold line bundle $\mathcal{L}^{-1}$ by contracting the zero section. For simplicity, we will call $o\in\mathcal{C}$ an orbifold cone singularity. The affine variety $\mathcal{C}$ can be compactified to a projective variety $\overline{\mathcal{C}}$ by adding an divisor $\mathcal{Y}_{\infty}\cong\mathcal{Y}$ at infinity (see [Kol04, 14]). According to [Kol04, 42], the singularity $o\in\mathcal{C}$ is klt if and only $(Y,\Delta)$ is a Fano orbifold (i.e. the orbifold anti-canonical line bundle $K^{\mathrm{orb}}_{(Y,\Delta)}$ is ample) and there exists $\ell\in\mathbb{Q}_{>0}$ such that $K^{\mathrm{orb}}_{(Y,\Delta)}\sim_{\mathbb{Q}}\ell\mathcal{L}$ (see Proposition 2.15 for a proof in our setting). In this case, we will simply call the orbifold cone $o\in\mathcal{C}$ a Fano cone (by following the terminology in [LLX20, Definition 3.7]). Moreover, in the quasi-regular case, the metrics $g_{0}$ and $g^{M}$ discussed in the previous paragraph are induced by a Hermitian metric $h$ on $\mathcal{L}$ over $(Y,\Delta)$ whose Chern curvature will be denoted by $\omega^{Y}$.
@@
Let $(W,g)$ be a complete Kähler manifold whose integrable complex structure is denoted by $J$. It is called asymptotically conical (AC) with the asymptotical cone $(\mathcal{C},g_{0})$ if there exists a compact subset $K\subset W$ and a diffeomorphism $\Phi:\{{\mathfrak{r}}>1\}\rightarrow W\setminus K$ such that for some $\lambda_{1}>0,\lambda_{2}>0$ and all $j\in\mathbb{Z}_{\geq 0}$.
@@
Example 2.6 (Tian-Yau metrics, [TY91]).
Let $X$ be a Fano manifold, which means that $X$ is a projective manifold with an ample anticanonical line bundle $K_{X}$. Let $Y$ be a smooth divisor whose associated line bundle over $X$ is denoted by $[Y]$. Assume that there exists $\beta>1$ such that $K_{X}=\beta[Y]$. In particular, $[Y]$ is an ample line bundle. By adjunction formula, $K_{Y}=(\beta-1)L$ with $L=[Y]|_{Y}$ being the normal bundle of $Y\subset X$ and hence $Y$ is also a Fano manifold. Then there exists a complete AC Kähler metric on $X\setminus Y$ whose asymptotical cone $\mathcal{C}$ is given by $\mathrm{Spec}\left(\bigoplus_{m=0}^{+\infty}H^{0}(Y,mL)\right)$. In fact, the Kähler form near $Y$ can be chosen to be given by $\sqrt{-1}\partial\bar{\partial}\|s\|^{-\frac{2(\beta-1)}{n}}$ where $s$ is the defining section of $Y$ and $\|\cdot\|^{2}$ is a smooth Hermitian metric on the line bundle $[Y]$ ([TY91, (2.2)], see also [Li20, section 5]). Moreover, if $Y$ admits a Kähler-Einstein metric with positive Ricci curvature, then there exists a complete AC Ricci-flat Kähler metric on $X\setminus Y$ by solving a complex Monge-Ampère equation.
We will use the following compactification result, which only assumes that the complex structure is asymptotic to the cone.
@@
Let $(W,g,J)$ be a complete Kähler manifold and $(\mathcal{C},g_{0},J_{0})$ be an orbifold cone as before. Assume that there exist a compact subset $K\subset W$ and a diffeomorphism $\Phi:\{\mathfrak{r}>1\}\rightarrow W\setminus K$ such that for some $\lambda_{2}>0$ and all $j\in\mathbb{Z}_{\geq 0}$,
@@
    ([Li20, CH24]) Assume that the Reeb vector field $\mathfrak{v}$} is quasi-regular on the cone $\mathcal{C}$. There exists a compact Kähler orbifold $\mathcal{X}$ and a sub-orbifold divisor $\mathcal{Y}$ such that $\mathcal{Y}\cong\mathcal{C}^{\circ}/\langle\mathfrak{v}\rangle$ and $\mathcal{X}\setminus\mathcal{Y}$ is biholomorphic to $W$. Moreover there exists an orbifold diffeomorphism $\phi:(\mathcal{U},\mathcal{Y})\rightarrow(\mathcal{V},\mathcal{Y}_{\infty})$ where $\mathcal{U}$ is a strongly pseudo-concave neighborhood of $\mathcal{Y}\subset\mathcal{X}$ and $\mathcal{V}$ is a strongly pseudo-concave neighborhood of $\mathcal{Y}_{\infty}\subset\overline{\mathcal{C}}$ where $\overline{\mathcal{C}}=\mathcal{C}\cup\mathcal{Y}$ is the projective compactification of $\mathcal{C}$ mentioned above.